OpenStudy (nathanjhw):

The differential equation dy/dx = y/x^2 has a solution given by:

OpenStudy (nathanjhw):

y= C sqrt(x^2+4) y= Ce^(-1/x) y=1/2e^(x^2) + C x^2+y^2=C y= Ce^(-kt)

OpenStudy (anonymous):

Hi!

OpenStudy (nathanjhw):

Hi

OpenStudy (anonymous):

1. Compute the integral Int(x*exp(-x^2),x=0..1) a) 1/e b) 1/2 c) (1/2)(1-1/e) d) (1/e-1) e) -1/(2e) SOLUTION: The answer is C. The antiderivative is 2 - 1/2 exp(-x ) Inserting the limits 0 and 1, we obtain -(1-1/e)/2. which is response C. 2. There is a chemical reaction involving the compound N_2 O_5. Let C(t) denote the concentration of N_2 O_5 as a function of time t measured in seconds. This concentration satisfies dC/dt = -0.05C(t) How long (in seconds) will the reaction take to reduce the concentration to 10% of its original value? a) 10/(e^(-.05)) b) (e^(.05))/10 c) .05/ln(10) d) .05 ln(1/10) e) ln(10)/(.05) SOLUTION: The answer is E. The solution to the given differential equation is C(t)=K e^(-.05t), where K is the inititial concentration. We seek the time t when C(t)=K/10 (one tenth the initital concentration). Setting K/10=K e^(-.05t) and solving for t (by taking natural logs) gives t=ln(1/10)/(-.05). Since ln(1/10)= -ln(10), this becomes t=ln(10)/(.05), which is response E.

OpenStudy (anonymous):

u can see in this example equation the form used

OpenStudy (anonymous):

This equation is separable. Get y and dy on the same side and x and dx on the other.

OpenStudy (anonymous):

this is how we will solve ours

OpenStudy (anonymous):

excuse me @peachpi do not interrupt please, it is rude.

OpenStudy (nathanjhw):

@pandaluvs I'm confused. What am I supposed to do?

OpenStudy (anonymous):

the answer is E

OpenStudy (nathanjhw):

How do you know that?

OpenStudy (anonymous):

done the equation before

OpenStudy (anonymous):

:D :)

OpenStudy (nathanjhw):

Alright thanks

OpenStudy (anonymous):

no prob.

OpenStudy (anonymous):

um, no. FYI \[\frac{ dy }{ dx }=\frac{ y }{ x^2 }\] Cross multiply to separate the variables \[\frac{ dy }{ y }=\frac{ dx}{ x^2 }\] Integrate both sides. Put the constant of integration on the right \[\int\limits\frac{ dy }{ y }=\int\limits\frac{ dx}{ x^2 }\] \[\ln y = {-\frac{ 1 }{ x }+C}\] Make both sides the exponent of e \[y=e ^{-\frac{ 1 }{ x }+C}\] Break apart the exponent \[y=e ^{-\frac{ 1 }{ x }}*e^C\] Rewrite e^C as one constant \[y=Ce ^{-\frac{ 1 }{ x }}\]

OpenStudy (nathanjhw):

That's the answer I got as well after looking over it again. I was not so sure about panda's answer....

OpenStudy (nathanjhw):

Thanks @peachpi